Which of the numbers 9, 10, or 11 is the solution of 83 = 93 - x?

Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

Which of the numbers 9, 10, or 11 is the solution of 83 = 93 - x?
----------
Replace the x by 9 then by 10 then by 11 to see which
number makes the equation true.
10 is the answer.
Cheers,
Stan H.

Answer by jim_thompson5910(35256)

83=93-9 Plug in x=9

83=84 Combine like terms. Since the equation is not true, x=9 is not a solution

--------------------------------------------------------------------------------


83=93-10 Plug in x=10

83=83 Combine like terms. Since the equation is true, x=10 is a solution

--------------------------------------------------------------------------------


83=93-11 Plug in x=11

83=82 Combine like terms. Since the equation is not true, x=11 is not a solution

--------------------------------------------------------------------------------


Answer:

So the only solution is x=10